OpenStudy (anonymous):

Qn: what is the unit digit in {((264)^102)+((264)^103)} ? ans: ((264)^102)+((264)^103)=((264)^102){1+264}=((264)^102)+265 required unit digit = unit digit in [(4)^102 * 5] = unit digit in [((4)^4)^25 + 4^2 * 5] = unit digit in (6*6*5)=0 I really cant get through this logic...how did this happen??

OpenStudy (anonymous):

HINT: Unit digit means taking the modulus 10

OpenStudy (anonymous):

Also, \( 264^{102}+264^{103}=264^{102}{(1+264)}= 264^{102} \times 265 \)

OpenStudy (ash2326):

It's written a little wrong \[264^{102}+264^{103}\] \[264^{102}(1+264)\] \[264^{102} \times 265\] Now if we have 14 , and we need to find the unit digit in 14^2=196, we know it's 4 but it can also be found out by finding the unit digit of 4^2=>16, it's also 6 so in place of 264, we'll use 4, and for 265 we'll use 5. Th unit digit will be the unit digit of the product \[4^{102} \times 5\] 4^{102} is an even no. Any even no. multiplied by 5, will have 0 as the unit digit. So the unit digit is 0

OpenStudy (anonymous):

\[ 264^{102} \equiv 4^{102} \equiv 6^{51} \equiv 6 \mod 10 \]

OpenStudy (ash2326):

we know it's "6" Sorry made a typographical error

OpenStudy (anonymous):

and \( 265 \equiv 5 \pmod {10} \) combining we get \( 264^{102} \times 265 \equiv 0 \pmod{10} \) Hence the answer.

OpenStudy (anonymous):

all this is right, but i maybe it is easier to see

OpenStudy (anonymous):

if we start with \[264^{102}(1+264)\] and then \[264^{102}\times 265\] then all we need to know is that \[264^{102}\] is even

OpenStudy (anonymous):

Lolz :D Indeed sat!

OpenStudy (anonymous):

not sure why all the other stuff is necessary

OpenStudy (anonymous):

Others stuffs are necessary because it's the formal way of doing things but this approach is a neat one :D